OpenStudy (anonymous):

Given the equation 5x+10= -30, solve for x and choose the correct solution.

OpenStudy (diyadiya):

First step is to subtract 10 from both sides

OpenStudy (anonymous):

5x= -30-10 5x = -40 x = -40/5 x= -8

OpenStudy (anonymous):

-8